# Train for the Worst, Plan for the Best: Understanding Token Ordering in Masked Diffusions

### v2

@@ Abstract
In recent years, masked diffusion models (MDMs) have emerged as a promising alternative approach for generative modeling over discrete domains. Compared to autoregressive models (ARMs), MDMs trade off complexity at training time with flexibility at inference time. At training time, they must learn to solve an exponentially large number of infilling problems, but at inference time, they can decode tokens in essentially arbitrary order. In this work, we closely examine these two competing effects. On the training front, we theoretically and empirically demonstrate that MDMs indeed train on computationally intractable subproblems compared to their autoregressive counterparts. On the inference front, we show that a suitable strategy for adaptively choosing the token decoding order significantly enhances the capabilities of MDMs, allowing them to sidestep hard subproblems. On logic puzzles like Sudoku, we show that adaptive inference can boost solving accuracy in pretrained MDMs from $<7$% to $\approx 90$%, even outperforming ARMs with 7×7$\times$7 × as many parameters and that were explicitly trained via teacher forcing to learn the right order of decoding.
@@ 1 Introduction
In many applications, such as language modeling, masked diffusion models (MDMs) still underperform compared to autoregressive models (ARMs) (Nie et al., 2024; Zheng et al., 2024), which instead learn to reverse a noise process that unmasks tokens sequentially from left to right. However, recent studies suggest that MDMs may offer advantages in areas where ARMs fall short, including reasoning (Nie et al., 2024; Kitouni et al., 2025), planning (Ye et al., 2024), and infilling (Gong et al., 2024). This raises a key question: what are the strengths and limitations of MDMs compared to ARMs, and under what conditions can MDMs be scaled to challenge the dominance of ARMs in discrete generative modeling?
@@
  • Complexity at training time: By design, the prediction task that MDMs are trained on is more challenging. Whereas ARMs seek to predict the next token given an unmasked prefix, MDMs seek to predict a token conditioned on a set of unmasked tokens in arbitrary positions.
  • Flexibility at inference time: On the other hand, the sampling paths taken by an MDM are less rigid. The order in which tokens are decoded at inference time is random instead of fixed to left-to-right. In fact, even more is possible: MDMs can actually be used to decode in any order (Zheng et al., 2024).
@@
We prove that even for simple, benign models of data, there are noise levels at which a large fraction, but not all, of the corresponding subproblems solved by MDMs are computationally intractable. We then show this imbalance in computational complexity across subproblems persists even in real-world text data (Fig. 2, left).
@@
While the above might appear to be bad news for MDMs, in the second part of this paper we answer our guiding question in the affirmative by building upon the observation (Zheng et al., 2024) that MDMs which can perfectly solve all masking subproblems can be used to decode in any order.
In place of vanilla MDM inference whereby tokens are unmasked in random order, we consider adaptive strategies that carefully select which token to unmask next. Our key insight is that this adaptivity makes it possible to sidestep the hard subproblems from training (Fig. 1). In fact, we find that even without modifying how MDMs are trained, the resulting models’ logits contain enough information to determine the right order in which to unmask.
Our main empirical result is to show that the performance of MDMs pretrained on logic puzzle data dramatically improves when one goes from vanilla to adaptive inference. For example, on Sudoku puzzles, a simple adaptive strategy (Section 4.1) improves the accuracy of MDMs from $<7$% to almost 90%. Remarkably, this not only outperforms vanilla ARMs, but even bespoke ARMs trained to learn the right decoding order via supervised teacher forcing (Shah et al., 2024; Lehnert et al., 2024) (Table 2).
@@ 2 Masked Diffusion Models (MDM)
In this section, we explain the framework of Masked Diffusion Models (Shi et al., 2024; Sahoo et al., 2025) and its interpretation as an order-agnostic learner. MDMs gradually add noise to the true discrete data and learn the marginal distribution of the induced reverse process. Below, we formulate the forward and reverse processes for MDMs.
Let the distribution $p_{\rm{data}}$ on $\{1,\ldots,m\}^{L}$ be the data distribution over sequences of length $L$ and with vocabulary $\{1,\ldots,m\}$. We use 00 to denote the “mask” token.
@@
$q_{t|0}(x_{t}^{i}\mid x_{0}^{i})=\mathrm{Cat}\bigl{(}\alpha_{t}\mathbf{e}_{x_{0}^{i}}+(1-\alpha_{t})\mathbf{e}_{0}\bigr{)}\,,$
where $\alpha_{t}$ is the predefined noise schedule satisfying $\alpha_{0}\approx 1,\alpha_{1}\approx 0$ and $e_{x_{0}^{i}}\in\mathbb{R}^{m+1}$ denotes a one-hot vector corresponding to the value of token $x_{0}^{i}$. $\mathrm{Cat}(\pi)$ denotes the categorical distribution given by $\pi\in\Delta^{m}$. In other words, for each $i$-th coordinate, $x_{t}^{i}$ is masked to the mask token 00 with probability $1-\alpha_{t}$ and unchanged otherwise.
@@
The reverse process of the above forward process is denoted using $q_{s|t}(x_{s}|x_{t},x_{0})$ and is given by $q_{s|t}(x_{s}|x_{t},x_{0})=\prod_{i=0}^{L-1}q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ for any $s<t$, where
$q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0})=\begin{cases}\mathrm{Cat}(\mathbf{e}_{x_{t}^{i}})\quad&x_{t}^{i}\neq m\\ \mathrm{Cat}\left(\frac{1-\alpha_{s}}{1-\alpha_{t}}\mathbf{e}_{m}+\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\mathbf{e}_{x_{0}}\right)\quad&x_{t}^{i}=m\,.\end{cases}$
The reverse transition probability $q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ is approximated using $g_{\theta}(x_{s}^{i}|x_{t})\triangleq q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0}\leftarrow p_{\theta}(x_{t},t))$ where $p_{\theta}(x_{t},t)$ is a denoising network trained to predict the marginal on $x_{0}$ via an ELBO-based loss. To be precise, $q_{s|t}\left(x_{s}^{i}\mid x_{t},x_{0}\leftarrow p_{\theta}(x_{t},t)\right)$ indicates the conditional probability where $p_{\theta}(x_{t},t)$ is placed in the position of $x_{0}$ within $q_{s|t}(x_{s}^{i}\mid x_{t},x_{0})$.
$\mathcal{L}_{\theta}=\int_{0}^{1}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}x_{0}\sim p_{\rm data}\\ x_{t}\sim q_{t|0}(\cdot|x_{0})\end{subarray}}\left[\delta_{x_{t},0}\mathbf{e}_{x_{0}}^{\intercal}\log p_{\theta}(x_{t},t)\right]dt.$
Here, $\alpha_{t}^{\prime}=\frac{d\alpha_{t}}{dt}$ and $\delta_{x_{t},0}$ is the indicator function; the summation is computed over coordinates $i$ s.t. $x_{t}^{i}=0$. In practice, a time-embedding-free architecture for the denoising network, i.e., $p_{\theta}(x_{t},t)=p_{\theta}(x_{t})$, is usually employed as $x_{t}$ implicitly contains information about $t$ via the number of masked tokens.
The reverse sampling process starts from the fully masked sentence $x_{1}=(0,\ldots,0)$. At a given noise level $t\in(0,1]$, suppose we have a partially masked sequence $x_{t}$. For predetermined noise level $s<t$, we sample $x_{s}\sim g_{\theta}(\cdot|x_{t})$. This process is repeated recursively from $t=1$ to $t=0$.
@@ 2.1 Reformulating the training and inference of MDMs
In this section, we first discuss vanilla order-agnostic training of MDMs and compare it with “left-to-right” order training of autoregressive models in Section 2.1.1. Then, we reformulate vanilla MDM inference in Section 2.1.2 to set the stage for the upcoming discussion.
@@ 2.1.1 Order-agnostic training of MDMs
Recent works (Zheng et al., 2024; Ou et al., 2024) have observed that the learning problem of MDM is equivalent to a masked language model. Building upon their analysis, we reformulate the loss $\mathcal{L}_{\theta}$ to show that $\mathcal{L}_{\theta}$ is a linear combination of the loss for all possible infilling masks. We first define $x_{0}[M]$ as a masked sequence, obtained from original sequence $x_{0}$ where indices in the mask set $M$ (regarded as a subset of $[L]\triangleq\{1,2,\ldots,L\}$) are replaced with mask token 00.
@@
$\mathcal{L}_{\theta}=-\frac{1}{L}\sum_{M\subseteq[L],i\in M}\frac{1}{\binom{L-1}{|M|-1}}\displaystyle\mathop{\mathbb{E}}_{x_{0}\sim p_{\rm data}}[\log p_{\theta}(x^{i}_{0}|x_{0}[M])],$ (1)
where $|M|$ denote the size of the mask $M$ and $p_{\theta}(x_{i}\mid x_{0}[M])$ indicates the conditional probability of the $i$-th coordinate from $p_{\theta}(x_{t})$.
The proof of the above proposition is given in Appendix E. As the MDM loss is a linear combination of the loss for all possible infilling mask $M$, the minimizer of the loss $\mathcal{L}_{\theta}$ learns to solve every masking problem. More formally, for all subsets $M\subseteq\{1,2,\ldots,L\}$, we have $\arg\min_{\theta}\log p_{\theta}(x^{i}_{0}|x_{0}[M])=p_{\rm{data}}(x^{i}_{0}|x_{0}[M])$. In other words, the optimal predictor $p_{\theta}$ is the posterior marginal of the $i$-th token, conditioned on $x_{0}[M]$ for all masks $M$. The training objective of MDM aims to predict $x_{0}$ from $x_{0}[M]$ across all possible masks. Hence, we will refer to the MDM training as order-agnostic training.
On the other hand, Autoregressive Models (ARMs) learn to solve a smaller set of infilling problems ($L$ infilling problems in ARMs as opposed to $\exp(L)$ infilling problems in MDM) by predicting $i^{\textrm{th}}$ token $x^{i}$ given all previous tokens $x^{0},\ldots,x^{i-1}$. This prediction problem is equivalent to predicting $x^{i}$ by masking at positions $\{i,\ldots,L-1\}$. Therefore, we can write it as
@@
ARMs are trained to predict tokens sequentially from left to right in all sequences. We refer to this as left-to-right training. In general, one can also consider predicting tokens sequentially under some fixed, known permutation of the sequence; we refer to this as order-aware training.
@@ 2.1.2 Order-agnostic inference of MDMs
[…]
  • (a)
    Sample a set of masked tokens $\mathcal{S}\subseteq\{i\mid x_{t}^{i}=0\}$, $\mathbb{P}(i\in\mathcal{S})=\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$.
  • (b)
    For each $i\in\mathcal{S}$, sample $x_{s}^{i}\sim p_{\theta}(x^{i}|x_{t})$.
@@ 3 MDMs train on hard problems
Figure 2: Left: MDMs train on hard problems (Section 3.2). x-axis and y-axis correspond to $\log(\text{FLOPs})$ and $-\log p_{\theta}(x)$, respectively. MDM (Blue) is worse than ARM (Orange) in likelihood modeling. Most masking problems (Other lines) that MDM is trained on are harder than those encountered by ARM, as indicated by small log-likelihoods. Right: Task error imbalance (Section 3.3). MDM’s performance varies across different tasks. For text data (top right), this is indicated by validation loss. For L&O-NAE-SAT (bottom right), MDM performs well on the masking problems for observation positions (light region) but struggles with latent positions (dark region).
In this section, we theoretically and empirically demonstrate that a large portion of masking subproblems $p_{\theta}(x^{i}_{0}\mid x_{0}[M])$ can be difficult to learn. For intuition, consider solving a masked prediction problem $p_{\theta}(x^{i}\mid x_{0}[M])$ on text data like masking an arbitrary sentence in the middle of a document and predicting the correct word for a specific position in that sentence. It is reasonable that this task should be more complex, even for humans, than left-to-right prediction, and in this section, we place this intuition on a rigorous footing.
In Section 3.1, we show several examples of simple, non-pathological distributions for which: (1) the masking problems encountered during order-aware training are computationally tractable, yet (2) many of the ones encountered during order-agnostic training are computationally intractable. In Section 3.2, we empirically show that text data also exhibits this gap between the computational complexity of order-aware and order-agnostic training. In Section 3.3, we reveal that this discrepancy in computational complexity manifests empirically in performance imbalance across tasks: as predicted by the theory, MDMs trained on data from such distributions exhibits small errors on easy subproblems but suffers from large errors on harder ones.

### v3

@@ Abstract
In recent years, masked diffusion models (MDMs) have emerged as a promising alternative approach for generative modeling over discrete domains. Compared to autoregressive models (ARMs), MDMs trade off complexity at training time with flexibility at inference time. At training time, they must learn to solve an exponentially large number of infilling problems, but at inference time, they can decode tokens in essentially arbitrary order. In this work, we closely examine these two competing effects. On the training front, we theoretically and empirically demonstrate that MDMs indeed train on computationally intractable subproblems compared to their autoregressive counterparts. On the inference front, we show that a suitable strategy for adaptively choosing the token decoding order significantly enhances the capabilities of MDMs, allowing them to sidestep hard subproblems. On logic puzzles like Sudoku, we show that adaptive inference can boost solving accuracy in pretrained MDMs from <7<7% to $\approx 90$%, even outperforming ARMs with $7\times$ as many parameters and that were explicitly trained via teacher forcing to learn the right order of decoding. This shows that MDMs without knowledge of the correct token generation order during training and inference can outperform ARMs trained with knowledge of the correct token generation order. We also show the effectiveness of adaptive MDM inference on reasoning tasks such as coding and math on the 8B large language diffusion model (LLaDa 8B).
@@ 1 Introduction
In many applications, such as language modeling, masked diffusion models (MDMs) still underperform compared to autoregressive models (ARMs) (Nie et al., 2024; Zheng et al., 2024), which instead learn to reverse a noise process that unmasks tokens sequentially from left to right. However, recent studies suggest that MDMs may offer advantages in areas where ARMs fall short, including reasoning (Nie et al., 2024; Kitouni et al., 2025), planning (Ye et al., 2024), and infilling (Gong et al., 2024). This raises a key question: what are the strengths and limitations of MDMs compared to ARMs, and on what type of tasks can MDMs be scaled to challenge the dominance of ARMs in discrete generative modeling?
@@
  • Complexity at training time: MDMs face a more challenging training task by design. While ARMs predict the next token given an unmasked prefix, MDMs predict a token conditioned on a set of unmasked tokens in arbitrary positions. This inherently increases their training complexity.
  • Flexibility at inference time: On the other hand, the sampling paths taken by an MDM are less rigid. Unlike the fixed left-to-right decoding of ARMs, MDMs decode tokens in random order at inference. Even more is possible: MDMs can be used to decode in any order (including left-to-right).
@@
Theoretically, we show examples of simple data distributions with a natural left-to-right order, where ARMs can provably generate samples efficiently. In contrast, there are noise levels at which a large fraction of the corresponding subproblems solved by MDMs for these distributions are provably computationally intractable. Empirically, we validate this claim on real-world text data, known to have left-to-right order and show that the imbalance in training complexity across subproblems persists even in real-world text data (Fig. 2, left).
@@
While the above might appear to be bad news for MDMs, in the second part of this paper, we answer our guiding question in the affirmative by building upon the observation (Chang et al., 2022; Zheng et al., 2023) that MDMs which can perfectly solve all masking subproblems can be used to decode in any order.
In first part of the paper, we show that the imbalance in complexity across subproblems during the training of MDMs results in some of the subproblems being poorly trained and the vanilla MDM inference that unmasks tokens in random order results in evaluating the poorly trained marginals. Therefore, in place of vanilla MDM inference, we consider adaptive strategies that carefully select which token to unmask next. Our key insight is that the adaptive strategies makes it possible to sidestep the hard subproblems from training (Fig. 1). In particular, we find that even without modifying how MDMs are trained, the resulting models’ logits contain enough information to determine the right order in which to unmask. We show the effectiveness of the adaptive inference in solving logic puzzles, coding, math and infilling tasks. For example, on Sudoku puzzles, a simple adaptive strategy (Section 4.1) improves the accuracy of MDMs from <7<7% to almost 90%.
Advantage of MDMs over ARMs.
We show that the main effectiveness of MDMs lies in tasks that do not have the same natural token generation order across all sequences (e.g., logic puzzles and reasoning tasks like coding and math). By carefully designing experiments on logic puzzles, we show that MDMs without the knowledge of the correct token generation order during training and inference can outperform ARMs trained with the knowledge of the correct token generation order. In particular, we show that MDMs that decide the correct token generation order during inference via adaptive strategies can outperform ARMs that are trained to learn the right token generation order via supervised teacher forcing (Shah et al., 2024; Lehnert et al., 2024).
@@ 2 Masked Diffusion Models (MDM)
In this section, we explain the framework of Masked Diffusion Models (Shi et al., 2024; Sahoo et al., 2025) and highlight its interpretation as an order-agnostic learner. MDMs gradually add masking noise to the true discrete data and learn the marginal distribution of the induced reverse process. We formally define both the forward and reverse processes for MDMs below.
Let the distribution $p_{\rm{data}}$ on $\{1,\ldots,m\}^{L}$ be the data distribution over sequences of length $L$ and with vocabulary $\{1,\ldots,m\}$. We use 0 to denote the “mask” token.
@@
$q_{t|0}(x_{t}^{i}\mid x_{0}^{i})=\mathrm{Cat}\bigl{(}\alpha_{t}\mathbf{e}_{x_{0}^{i}}+(1-\alpha_{t})\mathbf{e}_{0}\bigr{)}.$
Here, $\alpha_{t}$ is a predefined noise schedule satisfying $\alpha_{0}\approx 1,\alpha_{1}\approx 0$ and $\mathbf{e}_{x_{0}^{i}}\in\mathbb{R}^{m+1}$ is a one-hot vector corresponding to the value of token $x_{0}^{i}$. $\mathrm{Cat}(\pi)$ denotes the categorical distribution given by $\pi\in\Delta^{m}$. In other words, for each $i$-th coordinate, $x_{t}^{i}$ is masked to the mask token 0 with probability $1-\alpha_{t}$ and remains unchanged otherwise.
@@
The reverse process of the above forward process is denoted by $q_{s|t}(x_{s}|x_{t},x_{0})$ and is given by $q_{s|t}(x_{s}|x_{t},x_{0})=\prod_{i=0}^{L-1}q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ for any $s<t$, where
$q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0})=\begin{cases}\mathrm{Cat}(\mathbf{e}_{x_{t}^{i}})\quad&x_{t}^{i}\neq 0\\ \mathrm{Cat}\left(\frac{1-\alpha_{s}}{1-\alpha_{t}}\mathbf{e}_{0}+\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\mathbf{e}_{x_{0}^{i}}\right)\quad&x_{t}^{i}=0\,.\end{cases}$
The reverse transition probability $q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ is approximated using $g_{\theta}(x_{s}^{i}|x_{t})\triangleq q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0}\leftarrow p_{\theta}(\cdot|x_{t},t))$ where $p_{\theta}(\cdot|x_{t},t)$ is a denoising network trained to predict the marginal distribution on $x_{0}^{i}$ via an ELBO-based loss for all masked tokens at noise scale $t$ (i.e., for all $i$ such that $x_{t}^{i}=0$). To be precise, $q_{s|t}\left(x_{s}^{i}\mid x_{t},x_{0}\leftarrow p_{\theta}(\cdot|x_{t},t)\right)$ indicates the conditional probability where $p_{\theta}(\cdot|x_{t},t)$ is placed in the position of $e_{x_{0}^{i}}$ within $q_{s|t}(x_{s}^{i}\mid x_{t},x_{0})$. The denoising network is trained to minimize the following loss derived from the score-entropy (Lou et al., 2024; Sahoo et al., 2025; Shi et al., 2024; Ou et al., 2024):
$\mathcal{L}_{\theta}=\int_{0}^{1}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}x_{0}\sim p_{\rm data}\\ x_{t}\sim q_{t|0}(\cdot|x_{0})\end{subarray}}\sum_{i:x_{t}^{i}=0}-\log p_{\theta}(x_{0}^{i}|x_{t},t)dt,$
where $\alpha_{t}^{\prime}=\frac{d\alpha_{t}}{dt}$ and the summation is computed over masked tokens (i.e., all $i$ such that $x_{t}^{i}=0$). In practice, a time-embedding-free architecture for the denoising network, i.e., $p_{\theta}(\cdot|x_{t},t)=p_{\theta}(\cdot|x_{t})$ is generally used as $x_{t}$ implicitly contains information about $t$ via the number of masked tokens.
[…]
@@ 2.1 Reformulating the training and inference of MDMs
In this section, we first discuss training of MDMs and compare it with “left-to-right” order training of autoregressive models in Section 2.1.1. Then, we reformulate vanilla MDM inference in Section 2.1.2 to set the stage for the upcoming discussion.
@@ 2.1.1 Order-agnostic training of MDMs
Recent works (Zheng et al., 2024; Ou et al., 2024) have observed that the learning problem of MDM is equivalent to a masked language model. Building upon their analysis, we reformulate the loss $\mathcal{L}_{\theta}$ to show that $\mathcal{L}_{\theta}$ is a linear combination of the loss for all possible infilling masks. We first define $x_{0}[M]$ as a masked sequence, obtained from original sequence $x_{0}$ where indices in the mask set $M$ (a subset of $L]\triangleq\{1,2,\ldots,L\}$) are replaced with mask token 0.
@@
$\mathcal{L}_{\theta}=-\sum_{M\subseteq[L],i\in M}\frac{1}{|M|}\frac{1}{\binom{L}{|M|}}\displaystyle\mathop{\mathbb{E}}_{x_{0}\sim p_{\rm data}}[\log p_{\theta}(x^{i}_{0}|x_{0}[M])],$ (1)
where $|M|$ is the size of the set $M$ and $p_{\theta}(x_{i}\mid x_{0}[M])$ indicates the conditional probability of the $i$-th coordinate from $p_{\theta}(x_{t})$.
The proof of the above proposition is given in Appendix E. As the MDM loss is a linear combination of the loss for all possible infilling mask $M$, the minimizer of the loss $\mathcal{L}_{\theta}$ learns to solve every masking problem. In other words, the optimal predictor $p_{\theta}$ is the posterior marginal of the $i$-th token, conditioned on $x_{0}[M]$ for all masks $M$.
On the other hand, Autoregressive Models (ARMs) learn to predict $i^{\textrm{th}}$ token $x^{i}$ based on all preceding tokens, from $x^{0}$ to $x^{i-1}$. This is equivalent to predicting $x^{i}$ by masking positions from $i$ to L1L-1. Therefore, the training objective for ARMs can be expressed as:
@@
Typically, ARMs are trained to predict tokens sequentially from left to right. We refer to this as left-to-right training. However, it’s also possible to train these models to predict tokens sequentially based on a fixed, known permutation of the sequence. We refer to this general approach as order-aware training.
To understand the comparison between the training objective of MDMs and ARMs, we want to highlight the equivalence between any-order autoregressive loss and MDM loss (Hoogeboom et al., 2021a; Ou et al., 2024). In particular, under conditions of Proposition 2.1, MDM loss is equal to
$\mathcal{L}_{\theta}=-\mathop{\mathbb{E}}_{\begin{subarray}{c}x_{0}\sim p_{\textrm{data}}\\ \pi\sim\textrm{Unif}(\mathbb{S}_{L})\end{subarray}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right],$
where $\textrm{Unif}(\mathbb{S}_{L})$ is a uniform distribution over all the permutations of length $L$ (See Appendix E.1 for the proof). Observe that if the expectation is only with respect to the identity permutation, then the loss becomes an autoregressive loss. This shows that MDM loss solves exponentially more subproblems than ARM loss. In contrast to ARM loss, MDM does not prefer any particular (e.g., left-to-right) order during the training; therefore, we call its training order-agnostic training.
@@ 2.1.2 Order-agnostic inference of MDMs
(a) Sample a set of masked tokens $\mathcal{S}\subseteq\{i\mid x_{t}^{i}=0\}$, $\mathbb{P}(i\in\mathcal{S})=\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$. (b) For each $i\in\mathcal{S}$, sample $x_{s}^{i}\sim p_{\theta}(x^{i}|x_{t})$.
@@
On the other hand, ARMs are trained to predict tokens sequentially from left to right and therefore, generate tokens also in left-to-right order. In contrast, vanilla MDM inference generates the tokens in a random order.
@@ 3 MDMs train on hard problems
Figure 2: Left: MDMs train on hard problems (Section 3.2). x-axis and y-axis correspond to $\log(\text{FLOPs})$ and $\log p_{\theta}(x)$, respectively. MDM (Blue) is worse than ARM (Orange) in likelihood modeling. Most masking problems (Other lines) that MDM is trained on are harder than those encountered by ARM, as indicated by small log-likelihoods. Right: Task error imbalance (Section 3.3). MDM’s performance varies across different tasks. For text data (top right), this is indicated by validation loss. For L&O-NAE-SAT (bottom right), MDM performs well on the masking problems for observation positions (light region) but struggles with latent positions (dark region).
In this section, we provide theoretical and empirical evidence that when the data distribution has left-to-right order (or any fixed known order) then autoregressive training in left-to-right order (or in the known order) is more tractable than MDMs. In particular, for such distributions with fixed order, we show that ARMs can efficiently sample from the distributions but for MDMs, we theoretically and empirically demonstrate that a large portion of masking subproblems $p_{\theta}(x^{i}_{0}\mid x_{0}[M])$ can be difficult to learn.
In Section 3.1, we show several examples of simple, non-pathological distributions for which: (1) the masking problems encountered during order-aware training (such as in ARMs) are computationally tractable, yet (2) many of the ones encountered during order-agnostic training (such as in MDMs) are computationally intractable. In Section 3.2, we empirically show that text data also exhibits this gap between the computational complexity of order-aware and order-agnostic training and therefore, MDMs train on subproblems of wide variety of complexity (depending on the order/masks). In Section 3.3, we empirically show that the variety in training complexity results in performance imbalance across subproblems: MDMs trained on data from such distributions exhibits small errors on easy subproblems but suffers from large errors on harder ones.